OpenStudy (anonymous):

What is the solution of the equation? √2x+13-5=x

OpenStudy (aravindg):

Take all x terms to one side and constant to other side

OpenStudy (anonymous):

\[\sqrt{2x+13}-5=x\]??

OpenStudy (anonymous):

I got x=-6 and -2, is that right?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

-6 wont work

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i assume you squared at one point and then solved a quadratic equation right?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

when you square , you can introduce an extraneous solution just like you know \(-5\neq 5\) but \((-5)^2=5^2\)

OpenStudy (anonymous):

the point is that you have to check your answers against the original equation

OpenStudy (anonymous):

that's confusing,,

OpenStudy (anonymous):

\[\sqrt{2x+13}=x+5\] put \(x=-6\) get \[\sqrt{2\times (-6)=13}=-6+5\] \[\sqrt{1}=-1\] which is false

OpenStudy (anonymous):

the second?

OpenStudy (anonymous):

\(-2\) should work

OpenStudy (anonymous):

\[\sqrt{2\times(-2)+13}=-2+5\] \[\sqrt{9}=3\] which is true

OpenStudy (anonymous):

point is that when you square as one step, you have to check your solutions in the original equation to make sure they both work

OpenStudy (anonymous):

they both might work, or maybe one works and one does not

OpenStudy (anonymous):

oh okay so the answer is -2?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you!!